# Electric field induction in quark-gluon plasma due to thermoelectric effects

### v1

@@ 4.1.1 Without Landau quantization
$\bm{\sigma}=\begin{pmatrix}\sigma_{e}&-\sigma_{H}&\sigma_{H}\\ \sigma_{H}&\sigma_{e}&-\sigma_{H}\\ -\sigma_{H}&\sigma_{H}&\sigma_{e}\\ \end{pmatrix},~{}\hskip 17.07164pt\mathbb{E}=\begin{pmatrix}E_{x}\\ E_{y}\\ E_{z}\end{pmatrix},~{}$
$\mathbb{L}=\frac{1}{T}\begin{pmatrix}-L_{34}&L_{56}&-L_{56}\\ -L_{56}&-L_{34}&L_{56}\\ L_{56}&-L_{56}&-L_{34}\\ \end{pmatrix},~{}\mathbb{X}=\begin{pmatrix}\frac{dT}{dx}\vspace{2mm}\\ \frac{dT}{dy}\vspace{2mm}\\ \frac{dT}{dz}\\ \end{pmatrix}.$
@@
Now, in heavy ion collisions, the net magnetic field could possibly have a preferred direction Hongo:2013cqa . Therefore, for a special case when the magnetic field is directed along the y-axis (say), i.e., the magnetic field unit vector $h^{j}\equiv(0,1,0)$, Eq. (4.1.1) gives,
@@
$S_{B}$ $=\frac{(\frac{\sigma_{el}}{T})(\frac{L_{34}}{T^{2}})+(\frac{\sigma_{H}}{T})(\frac{L_{56}}{T^{2}})}{(\frac{\sigma_{el}}{T})^{2}+(\frac{\sigma_{H}}{T})^{2}},~{}~{}~{}\overline{S}_{B}=\frac{(\frac{L_{34}}{T^{2}})}{(\frac{\sigma_{el}}{T})},~{}~{}~{}NB=\frac{(\frac{\sigma_{el}}{T})(\frac{L_{56}}{T^{2}})-(\frac{\sigma_{H}}{T})(\frac{L_{34}}{T^{2}})}{(\frac{\sigma_{el}}{T})^{2}+(\frac{\sigma_{H}}{T})^{2}}~{}.$ (52)
Note that, at vanishing Hall conductivity ($\sigma_{H}=0$) the magneto-Seebeck coefficient ${S_{B}}$ becomes the same as $\overline{S}_{B}$. This can be realised as follows. The coefficient $\overline{S}_{B}$ is responsible for the electric field in the y-axis or along the magnetic field, where the Lorentz force has null contribution. Therefore, the expression of $\overline{S}_{B}$ is similar to that of $eB_{0}=0$ case as in Eq. (27), except the fact that the integral in $L_{34}$ depends on magnetic decay parameter $\tau_{B}$ at finite magnetic field. Furthermore, for a static picture, where cooling effects are absent $\left(\frac{dT}{dt}=0\right)$, the integrals $L_{34}$ and $L_{56}$ reduce to $L_{34}$ = $\sum_{i}q_{i}L_{3i}$ and $L_{56}$ = $\sum_{i}q_{i}L_{5i}$, then the expression of Seebeck and Nernst coefficients becomes similar to that obtained in earlier studies PhysRevD.102.096011 ; Das2021 .
4.1.2 With Landau quantization
Now, we study the effect of Landau quantization on the thermoelectric coefficient of the QGP medium and see how it deviates from classical results. Basic modifications will occur in the dispersion relation for medium constituents and the phase space integration. As we considered the magnetic field along the y direction therefore, momentum quantization will occur in its perpendicular plane, i.e., the x-z plane. Hence,
@@ 4.1.2 With Landau quantization
[…]
The factor 2 of spin degeneracy in left hand side of last line will be converted to $\alpha_{l}$, which will be 2 for all Landau levels $l$, except lowest Landau level (LLL) $l=0$, where $\alpha_{l}=1$. In general, one can write $\alpha_{l}=2-\delta_{l,0}$. Under the conditions of an extremely high magnetic field for which all medium constituents occupy the lowest Landau energy level $l=0$. It means that the perpendicular motion of medium constituents completely vanishes as $k_{x}\approx k_{z}\approx 0$ at $l=0$. However, below that strong magnetic field limt, $l>0$ energy levels might have some non-negligible contributions, and the lowest Landau level (LLL) approximation is not sufficient enough PhysRevC.106.044914 ; Dey:2021fbo . This is the case in our study, where the magnetic field is not constant but decays with time. Here, we also consider, $k_{x}^{2}\approx k_{z}^{2}\approx(\frac{k_{x}^{2}+k_{z}^{2}}{2})=\frac{2l|{q_{i}}|B}{2}$, the integrals in Eq. (4.1.1) can be expressed as,
@@
4.1.3 Analytical solution for massless non-interacting system
To find the analytical expressions for the massless non-interacting system, we follow the same procedure mentioned in Sec. (3.1.1). Hence, for the case of a massless non-interacting system, the integral in Eq. (4.1.1) turns into the following forms:
@@ 4.1.3 Analytical solution for massless non-interacting system
Here, the integrals $L_{34}$ and $L_{56}$ in Eq. (52) are proportional to $q_{i}$, hence for the massless non-interacting system $L_{34}$ and $L_{56}$ both individually summed upto zero. Therefore, dimensionless transport coefficients - two magneto-Seebeck coefficients ($S_{B}$, $\overline{S}_{B}$) and one Nernst coefficient ($NB$) in Eq. (52) vanishes for the massless non-interacting system, hence
@@ 4.1.4 Numerical solution for QGP EoS
Figure 4: Magneto-Seebeck coefficient ($S_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $S_{B}$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $S_{B}$ in Gubser flow for static and evolving case at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
Fig. (4) represents the magneto-Seebeck coefficient ($S_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV in the presence of time-varying magnetic field with initial value $eB_{0}=1~{}m_{\pi}^{2}$ (a) and (b) and $eB_{0}$ = 5 $m_{\pi}^{2}$ (c) and (d) with its decay parameter $\tau_{B}$ = 3 fm, 7 fm. In all the plots, we observe similar behavior of $S_{B}$. For the case of Bjorken flow, at the early times, the coefficient $S_{B}$ is negative, which means that the direction of the produced electric field is opposite to the temperature gradient of the medium. Later, for a particular value of $\tau$, the $S_{B}$ becomes zero and then increases with positive value as $\tau$ increases. Here, we take the static and evolving pictures for both the cases along with Landau quantization (LQ) and without Landau quantization (w/o LQ). For the Bjorken flow case, the ideal MHD cooling rate is used as mentioned in Eq. (2.1.2). The black solid line represents the static picture without Landau quantization, and the green dash-dot-dash line represents the static picture with Landau quantization. On the other hand, the blue dash-dot line represents the evolving picture w/o LQ, and the red dashed line represents the evolving picture with LQ. $S_{B}$ is highly sensitive to magnetic decay parameter $\tau_{B}$. In the lower panels of (a) and (c), the percentage deviation without LQ from LQ is plotted with a magenta dash-dotted line. As expected, the effect of LQ improves with the strength of the magnetic field ($eB_{0}$). The cyan dashed line represents the percentage deviation from static to evolving case. In both plots, this deviation is nearly 40%-45% in the early evolution and decreases as the medium evolves. As the value of $S_{B}$ becomes positive, this deviation shows a peak by shifting from a negative value to a positive value. Once this percentage deviation achieves its maximum value, later on, it starts to decrease with proper time. However, the validity of the peak-like structure in the deviation plots is questionable. Because as we mentioned before, the temperature falls beyond $0.150$ GeV at the corresponding times, where the quasiparticle model is not justified. In Fig. (4) (b) and (d), the upper and bottom panels correspond to decay parameter $\tau_{B}$ = 3 fm and $\tau_{B}$ = 7 fm, respectively. Here, we use the cooling rate obtained from Gubser flow as mentioned in Eq. (13). Here, we also study the effect of radial expansion of medium on coefficient $S_{B}$. The red solid line corresponds to $r=0$ fm, and the green dashed line corresponds to $r=3$ fm in the static picture. On the other hand, The blue dash-triple dot line corresponds to $r=0$ fm, and the black dashed line corresponds to $r=3$ fm in the evolving picture. $S_{B}$ follows the similar trend in both the plots. The numerical value of $S_{B}$ increases as a function of $\tau$. The effect of $\tau_{B}$ is identical as observed in the Bjorken flow case.
@@
Figure 5: Magneto-Seebeck coefficient ($\overline{S}_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $\overline{S}_{B}$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $\overline{S}_{B}$ in Gubser flow for static and evolving case at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
Fig. (5) is the same as Fig. (4), but for $\overline{S}_{B}$. From Eq. (52), at vanishing Hall like components $\sigma_{H}\rightarrow 0$ and $I_{56}\rightarrow 0$, $\overline{S}_{B}=S_{B}$. Expression of $\overline{S}_{B}$ is similar to that of Seebeck $S$ as in $eB=0$ case, except the fact that in magnetic field effective relaxation time modifies which is a function of $\tau_{R}$ and $\tau_{B}$. The Hall-like components ($\sigma_{H},I_{56}$) are approximately one order smaller than their leading order counterparts ($\sigma_{e},I_{34}$). Therefore, at the early evolution, $\overline{S}_{B}$ is close to $S_{B}$. In the later time, as the magnetic field significantly reduces, $\overline{S}_{B}$ approaches the Seebeck coefficient ($S$) as for the $eB=0$ case.
@@
Figure 6: Dimensionless Nernst coefficient ($NB$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $NB$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $NB$ in Gubser flow for static and evolving case at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
4.2 Induced electric field in anisotropic system
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 7: Electric field induced in the QGP in peripheral collisions with $eB_{0}=5~{}m_{\pi}^{2}$, $\tau_{B}=3$ fm.
Refer to captionRefer to caption
Figure 8: Magnitude of electric field in the transverse plane with $eB_{0}=5~{}m_{\pi}^{2}$.
Refer to captionRefer to caption
@@ 4.2 Induced electric field in anisotropic system
In Fig. (7), we plotted the space-time profile of electric field $\vec{E}$, induced in QGP created in peripheral collision with the initial magnetic field of $eB_{0}=5~{}m_{\pi}^{2}$. Upper panels are for magnetic field decay parameter $\tau_{B}=3$ fm at two different times $\tau=0.5,~{}3$ fm, and lower panels are for $\tau_{B}=7$ fm. Unlike the head-on collision case, $E_{z}\neq 0$ in the peripheral collision. Due to boost invariant symmetry, $\vec{E}$ is symmetric along the $z$ axis. However, symmetry in the transverse plane is now broken due to the magnetic field along the $y$ axis. This can be easily seen from the magnitude plot in Fig. (8). In the early time during the QGP evolution, the electric field could go as high as $eE\approx 1~{}m_{\pi}^{2}$, close to the QCD scale. We see that increasing $\tau_{B}$ from $3$ to $7$ fm, the strength of $\vec{E}$ reduces at an early time; however, at a later time, it is quite the same for $\tau_{B}=3$ and $7$ fm. This signifies that in the current formalism, the electric field induced by the thermoelectric effect increases for faster decay of the initially produced magnetic field. This can be understood as follows. For QGP with lower electrical conductivity, the decay of the magnetic field is comparatively fast. Now, from Eq. (52) thermoelectric coefficients varies inversely with electrical conductivity ($S_{B}/~{}\bar{S}_{B}/~{}NB\sim\frac{1}{\sigma_{e}}$). Therefore, smaller $\sigma_{e}$ means fast decay of magnetic field or smaller value of $\tau_{B}$, which enhances Seebeck and Nernet coefficients. As a result, $\vec{E}$ increases. In Fig. (9), we plotted $\vec{E}$ for $eB_{0}=1~{}m_{\pi}^{2}$ to see the effect of different magnetic field strengths. There is a minor change in the induced field upon changing the initial magnetic field strength from $eB=5$ to $1~{}m_{\pi}^{2}$. A similar thing was already observed in the thermoelectric coefficients in Figs. (45), and 6.

### v2

@@ IV.1.1 Without Landau quantization
$\bm{\sigma}=\begin{pmatrix}\sigma_{e}&-\sigma_{H}&\sigma_{H}\\ \sigma_{H}&\sigma_{e}&-\sigma_{H}\\ -\sigma_{H}&\sigma_{H}&\sigma_{e}\\ \end{pmatrix},~{}\hskip 17.07164pt\mathbf{E}=\begin{pmatrix}E_{x}\\ E_{y}\\ E_{z}\end{pmatrix},~{}$
$\mathbf{L}=\frac{1}{T}\begin{pmatrix}-L_{34}&L_{56}&-L_{56}\\ -L_{56}&-L_{34}&L_{56}\\ L_{56}&-L_{56}&-L_{34}\\ \end{pmatrix},~{}\mathbf{X}=\begin{pmatrix}\frac{dT}{dx}\vspace{2mm}\\ \frac{dT}{dy}\vspace{2mm}\\ \frac{dT}{dz}\\ \end{pmatrix}.$
@@
Now, in heavy-ion collisions, the net magnetic field could possibly have a preferred direction [55]. Therefore, for a special case when the magnetic field is directed along the y-axis (say), i.e., the magnetic field unit vector $h^{j}\equiv(0,1,0)$, Eq. (IV.1.1) gives,
@@
$S_{B}$ $=\frac{(\frac{\sigma_{el}}{T})(\frac{L_{34}}{T^{2}})+(\frac{\sigma_{H}}{T})(\frac{L_{56}}{T^{2}})}{(\frac{\sigma_{el}}{T})^{2}+(\frac{\sigma_{H}}{T})^{2}},~{}~{}~{}\overline{S}_{B}=\frac{(\frac{L_{34}}{T^{2}})}{(\frac{\sigma_{el}}{T})},$
$NB$ $=\frac{(\frac{\sigma_{el}}{T})(\frac{L_{56}}{T^{2}})-(\frac{\sigma_{H}}{T})(\frac{L_{34}}{T^{2}})}{(\frac{\sigma_{el}}{T})^{2}+(\frac{\sigma_{H}}{T})^{2}}~{}.$ (52)
Note that, at vanishing Hall conductivity ($\sigma_{H}=0$) the magneto-Seebeck coefficient ${S_{B}}$ becomes the same as $\overline{S}_{B}$. This can be realized as follows. The coefficient $\overline{S}_{B}$ is responsible for the electric field in the y-axis or along the magnetic field, where the Lorentz force has a null contribution. Therefore, the expression of $\overline{S}_{B}$ is similar to that of $eB_{0}=0$ case as in Eq. (27), except the fact that the integral in $L_{34}$ depends on magnetic decay parameter $\tau_{B}$ at finite magnetic field. Furthermore, for a static picture, where cooling effects are absent $\left(\frac{dT}{dt}=0\right)$, the integrals $L_{34}$ and $L_{56}$ reduce to $L_{34}$ = $\sum_{i}q_{i}L_{3i}$ and $L_{56}$ = $\sum_{i}q_{i}L_{5i}$, then the expression of Seebeck and Nernst coefficients becomes similar to that obtained in earlier studies [39, 38].
IV.1.2 With Landau quantization
Now, we study the effect of Landau quantization on the thermoelectric coefficients of the QGP medium and see how it deviates from classical results. Basic modifications will occur in the dispersion relation for medium constituents and the phase space integration. As we considered the magnetic field along the y direction therefore, momentum quantization will occur in its perpendicular plane, i.e., the x-z plane. Hence,
$\omega=(\vec{k}^{2}+m^{2})^{1/2}$$\rightarrow$ $\omega_{l}=(k_{y}^{2}+m^{2}+2l|{q_{i}}|B)^{1/2},$
@@ IV.1.2 With Landau quantization
The factor 2 of spin degeneracy in left hand side of last line will be converted to $\alpha_{l}$, which will be 2 for all Landau levels $l$, except lowest Landau level (LLL) $l=0$, where $\alpha_{l}=1$. In general, one can write $\alpha_{l}=2-\delta_{l,0}$. Under the conditions of an extremely high magnetic field for which all medium constituents occupy the lowest Landau energy level $l=0$. It means that the perpendicular motion of medium constituents completely vanishes as $k_{x}\approx k_{z}\approx 0$ at $l=0$. However, below that strong magnetic field limt, $l>0$ energy levels might have some non-negligible contributions, and the lowest Landau level (LLL) approximation is not sufficient enough [68, 35]. This is the case in our study, where the magnetic field is not constant but decays with time. Here, we also consider, $k_{x}^{2}\approx k_{z}^{2}\approx(\frac{k_{x}^{2}+k_{z}^{2}}{2})=\frac{2l|{q_{i}}|B}{2}$, the integrals in Eq. (IV.1.1) can be expressed as,
@@
IV.1.3 Analytical solution for massless noninteracting system
To find the analytical expressions for the massless noninteracting system, we follow the same procedure mentioned in Sec. (III.1.1). Hence, for the case of a massless noninteracting system, the integral in Eq. (IV.1.1) turns into the following forms:
@@ IV.1.3 Analytical solution for massless noninteracting system
Here, the integrals $L_{34}$ and $L_{56}$ in Eq. (52) are proportional to $q_{i}$, hence for the massless noninteracting system $L_{34}$ and $L_{56}$ both individually summed upto zero. Therefore, dimensionless transport coefficients - two magneto-Seebeck coefficients ($S_{B}$, $\overline{S}_{B}$) and one Nernst coefficient ($NB$) in Eq. (52) vanishes for the massless noninteracting system, hence
@@ IV.1.4 Numerical solution for QGP EoS
Figure 4: Magneto-Seebeck coefficient ($S_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $S_{B}$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $S_{B}$ in Gubser flow for static and evolving picture at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
Figure  (4) represents the magneto-Seebeck coefficient ($S_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV in the presence of time-varying magnetic field with initial value $eB_{0}=1~{}m_{\pi}^{2}$ (a) and (b) and $eB_{0}$ = 5 $m_{\pi}^{2}$ (c) and (d) with its decay parameter $\tau_{B}$ = 3 fm, 7 fm. In all the plots, we observe similar behavior of $S_{B}$. For the case of Bjorken flow, at the early times, the coefficient $S_{B}$ is negative, which means that the direction of the produced electric field is opposite to the temperature gradient of the medium. Later, for a particular value of $\tau$, the $S_{B}$ becomes zero and then increases with positive value as $\tau$ increases. Here, we take the static and evolving pictures for both cases along with Landau quantization (LQ) and without Landau quantization (w/o LQ). For the Bjorken flow case, the ideal MHD cooling rate is used as mentioned in Eq. (II.1.2). The black solid line represents the static picture without Landau quantization, and the green dash-dot-dash line represents the static picture with Landau quantization. On the other hand, the blue dash-dot line represents the evolving picture w/o LQ, and the red dashed line represents the evolving picture with LQ. $S_{B}$ is highly sensitive to magnetic decay parameter $\tau_{B}$. In the lower panels of (a) and (c), the percentage deviation without LQ from LQ is plotted with a magenta dash-dotted line. As expected, the effect of LQ improves with the strength of the magnetic field ($eB_{0}$). The cyan dashed line represents the percentage deviation from static to evolving picture. In both plots, this deviation is nearly 30%-35% in the early evolution and decreases as the medium evolves. As the value of $S_{B}$ becomes positive, this deviation shows a peak by shifting from a negative value to a positive value. Once this percentage deviation achieves its maximum value, later on, it starts to decrease with proper time. In Fig. (4) (b) and (d), the upper and bottom panels correspond to decay parameter $\tau_{B}$ = 3 fm and $\tau_{B}$ = 7 fm, respectively. Here, we use the cooling rate obtained from Gubser flow as mentioned in Eq. (II.2). Here, we also study the effect of radial expansion of medium on coefficient $S_{B}$. The red solid line corresponds to $r=0$ fm, and the green dashed line corresponds to $r=3$ fm in the static picture. On the other hand, The blue dash-triple dot line corresponds to $r=0$ fm, and the black dashed line corresponds to $r=3$ fm in the evolving picture. $S_{B}$ follows the similar trend in both the plots. The numerical value of $S_{B}$ increases as a function of $\tau$. The effect of $\tau_{B}$ is identical as observed in the Bjorken flow case.
@@
Figure 5: Magneto-Seebeck coefficient ($\overline{S}_{B}$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $\overline{S}_{B}$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $\overline{S}_{B}$ in Gubser flow for static and evolving picture at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
Figure (5) is the same as Fig. (4), but for $\overline{S}_{B}$. From Eq. (52), at vanishing Hall like components $\sigma_{H}\rightarrow 0$ and $I_{56}\rightarrow 0$, $\overline{S}_{B}=S_{B}$. Expression of $\overline{S}_{B}$ is similar to that of Seebeck $S$ as in $eB_{0}=0$ case, except the fact that in magnetic field effective relaxation time modifies which is a function of $\tau_{R}$ and $\tau_{B}$. The Hall-like components ($\sigma_{H},I_{56}$) are approximately one order smaller than their leading order counterparts ($\sigma_{e},I_{34}$). Therefore, at the early evolution, $\overline{S}_{B}$ is close to $S_{B}$. In the later time, as the magnetic field significantly reduces, $\overline{S}_{B}$ approaches the Seebeck coefficient ($S$) as for the $eB_{0}=0$ case.
@@
Figure 6: Dimensionless Nernst coefficient ($NB$) as a function of proper time ($\tau$) at $\mu_{B}=0.3$ GeV. (a) $NB$ for static, evolving, with and without (w/o) Landau quantization (LQ), four cases are plotted here in the upper panel for $\tau_{B}=3,~{}7$ fm for Bjorken flow at $1~{}m_{\pi}^{2}$. The lower panel represents their percentage deviation at $\tau_{B}=7~{}$fm. (b) $NB$ in Gubser flow for static and evolving picture at $r=0$ and $3$ fm with LQ is plotted at $\tau_{B}=3~{}$fm (upper panel) and $\tau_{B}=7~{}$fm (lower panel) at $eB_{0}=1~{}m_{\pi}^{2}$. (c) and (d) are the same as (a) and (b), respectively, but at $eB_{0}=5~{}m_{\pi}^{2}$.
@@
IV.2 Induced electric field in anisotropic system
Figure 7: Time evolution of the induced electric field in the QGP in the peripheral collisions with $eB_{0}=5~{}m_{\pi}^{2}$. Upper left: at = $\tau=0.5~{}$fm, $\tau_{B}=3~{}$fm, upper right: at $\tau=3~{}$fm, $\tau_{B}=3~{}$fm, Bottom left: at $\tau=0.5~{}$fm, $\tau_{B}=7~{}$fm, and bottom right: $\tau=3~{}$fm, $\tau_{B}=7~{}$fm.
Figure 8: Magnitude of the electric field in the transverse plane with $eB_{0}=5~{}m_{\pi}^{2}$.
@@ IV.2 Induced electric field in anisotropic system
In Fig. (7), we plotted the space-time profile of electric field $\vec{E}$, induced in QGP created in peripheral collision with the initial magnetic field of $eB_{0}=5~{}m_{\pi}^{2}$. Upper panels are for magnetic field decay parameter $\tau_{B}=3$ fm at two different times $\tau=0.5,~{}3$ fm, and lower panels are for $\tau_{B}=7$ fm. In the bottom left panel at $\tau=0.5$ fm and $\tau_{B}=7$ fm, the produced electric field direction differs from other results. Which needs a careful observation to understand. From the Eq. 51, components of electric field $E_{x},~{}E_{y},~{}E_{z}$ are directly proportional to $S_{B},~{}\bar{S}_{B},~{}NB$, respectively. Therefore, the relative strength of the thermoelectric coefficients is sensitive to the direction of the induced electric field. For $\tau_{B}=7$ fm at $\tau=0.5$ fm, $S_{B}$ and $\bar{S}_{B}$ dominates over $-NB$. As a result, $x$ and $y$ components of the electric field dominate. However, in other electric field results, $NB$ dominates, leading to electric field directing along the $z$ axis. Unlike the head-on collision case, $E_{z}\neq 0$ in the peripheral collision. Due to boost invariant symmetry, $\vec{E}$ is symmetric along the $z$ axis. However, symmetry in the transverse plane is now broken due to the magnetic field along the $y$ axis. This can be easily seen from the magnitude plot in Fig. (8). In the early time during the QGP evolution, the electric field could go as high as $eE\approx 1~{}m_{\pi}^{2}$, close to the QCD scale. We see that increasing $\tau_{B}$ from $3$ to $7$ fm, the strength of $\vec{E}$ reduces at an early time; however, at a later time, it is quite the same for $\tau_{B}=3$ and $7$ fm. This signifies that in the current formalism, the electric field induced by the thermoelectric effect increases for faster decay of the initially produced magnetic field. This can be understood as follows. For QGP with lower electrical conductivity, the decay of the magnetic field is comparatively fast. Now, from Eq. (52) thermoelectric coefficients varies inversely with electrical conductivity ($S_{B}/~{}\bar{S}_{B}/~{}NB\sim\frac{1}{\sigma_{e}}$). Therefore, smaller $\sigma_{e}$ means fast decay of magnetic field or smaller value of $\tau_{B}$, which enhances Seebeck and Nernst coefficients. As a result, $\vec{E}$ increases. In Fig. (9), we plotted $\vec{E}$ for $eB_{0}=1~{}m_{\pi}^{2}$ to see the effect of different magnetic field strengths. There is a minor change in the induced field upon changing the initial magnetic field strength from $eB=5$ to $1~{}m_{\pi}^{2}$. A similar thing was already observed in the thermoelectric coefficients in Figs. (45), and 6.
IV.2.1 Significance of the findings
Studies [69, 70, 71] show that charge-dependent directed flow ($v_{1}$) can be used as a probe to the electromagnetic field in the QGP. Recent experimental study at RHIC [22] has explored this with $v_{1}$ of $p^{\pm}$ (proton), $k^{\pm}$ (kaon), and $\pi^{\pm}$ (pion). In peripheral collisions, $v_{1}$ for positive (negative) charged particles get enhanced (reduced) in rapidity $y>0$, due to Hall effect from Lorentz force. This is reversed in rapidity $y<0$. Experimental data [22] show that $v_{1}$ for positively charged particles ($p^{+},~{}k^{+},~{}\pi^{+}$) dominates over their negative counterparts (or antiparticles). These findings support our results of the induced electric field. From Fig. 7, electric field in $x>0$ is directed along $+z$. Whereas, for $x<0$ it is directed along $-z$. This induced electric field enhances (reduces) the Hall effect in positively (negatively) charged particles by increasing (decreasing) their velocity from Coulombic force (not from spectators). As a result, $v_{1}$ for positively charged particles is higher than their antiparticles (see Fig. 5 of Ref. [22]). In future studies, we will extend this work to include a detailed quantitative analysis of the thermoelectric effect’s impact on $v_{1}$.